OpenStudy (henrietepurina):

The graph of f(x) = x2 is shifted 3 units to the left to obtain the graph of g(x). Which of the following equations best describes g(x)?

OpenStudy (henrietepurina):

@jim_thompson5910

OpenStudy (henrietepurina):

by x2 I meant x^2

OpenStudy (anonymous):

to shift to the left, change \[f(x)\] in to \(f(x+3)\)

OpenStudy (henrietepurina):

why +3 if we are going left, aka minus?

OpenStudy (anonymous):

the "it" is the graph, not the input

OpenStudy (anonymous):

what is really happening is the y axis is shifting 3 units to the right, which makes the graph appear to shift 3 units to the left lets do an example

OpenStudy (henrietepurina):

ok

OpenStudy (usukidoll):

I remember this... precalc stuff... but yeah anything in the parenthesis indicates left or right. without it you will be going up or down. Like for example f(x) = x^2+3 I am going up 3 units... f(x)=x^2-3 I'm going down 3 units

OpenStudy (anonymous):

we can even do it with \(f(x)=x^2\) that graph has the following ordered pairs \[\{-4,16),(-3,9),(-2,4),(-1,1),(0,0),(1,1),(2,4)\] and so on now lets take a look at the ordered pairs for \(f(x+3)=(x+3)^2\) and plug on those same numbers

OpenStudy (usukidoll):

f(x) = (x^2-3) I'm going 3 units to the right. f(x) =(x^2+3) I'm going 3 units to the left so what you do is you graph your x^2 and then shift it according to the directions.

OpenStudy (anonymous):

\[(-4,1),(-3,0),(-2,1),(-1,4),(0,9), (1,16)\] and so on `

OpenStudy (anonymous):

you are adding the 3 FIRST then evaluation the function it has the effect or translating the graph to the LEFT 3 units

OpenStudy (henrietepurina):

ok ok... im following so far... could the answer possibly be g(x) = (x + 3)^2?

OpenStudy (anonymous):

yes

OpenStudy (henrietepurina):

thanks for helping both of you, I immediatly understood what to do after your help :D

OpenStudy (anonymous):

yw